OpenStudy (anonymous):

How can I calculate the time taken for an element to decay by a certain amount?

OpenStudy (anonymous):

I have the decay constant and need to find out how long it will take for the element to decay from 5ug to 1ug

OpenStudy (anonymous):

Do you know differential equations or calculus?

OpenStudy (anonymous):

Yes, but I don't think this question requires it

OpenStudy (anonymous):

The question I'm tryingt o answer is: TA sample of 24Na has a half-life of 234 hours, How much time elapses before a 5μg sample contains 1μg of undecayed atoms?

OpenStudy (anonymous):

A sample of*

OpenStudy (anonymous):

\[\frac{1}{c}e^{-t \lambda}=ke^{-t \lambda}=N\]

OpenStudy (anonymous):

but I want to know how to do it myself

OpenStudy (anonymous):

whats c?

OpenStudy (anonymous):

Sorry, it deleted half of my answer

OpenStudy (anonymous):

lol, ok

OpenStudy (anonymous):

I have used a similar equation \[N = N _{0}e ^{-\lambda t}\]

OpenStudy (anonymous):

...in the past

OpenStudy (anonymous):

http://upload.wikimedia.org/math/a/8/a/a8a0bd0a12874474352b307d9e919076.png \[\frac{-1}{\lambda}(lnN+lnk)=\frac{-1}{\lambda}(\ln(kN))=t\] \[(\ln(kN))=-\lambda t\] \[kN=e^{-\lambda t}\] \[N=ce^{-\lambda t}\]

OpenStudy (anonymous):

Set c= sample at t=0 That is, \[ c= N_0 \]

OpenStudy (anonymous):

That's the derivation of it

OpenStudy (anonymous):

ah, so its the same as my equation?

OpenStudy (anonymous):

You're given the half life, so I'd convert it into 2^... instead of e^....

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I don't know how to get the values of N and N0 though, because that is the number of atoms, not the mass

OpenStudy (anonymous):

hendce why I haven't already used that equation

OpenStudy (anonymous):

\[N=N_0 (2^{\log_2(e)})^{-\lambda t}=N_02^{-\log_2(e) \lambda t}\]

OpenStudy (anonymous):

N_0= 5μg Just call it '5'- It doesn't matter which units you use (I suppose moles or no. of atoms would be most natural, though) as long as you're consistent

OpenStudy (anonymous):

but N refers to atoms, not mass (which is the 5 and the 1)

OpenStudy (anonymous):

So obviously the half-life \[= \log_2(e) \lambda\]

OpenStudy (anonymous):

Which element is it?

OpenStudy (anonymous):

Na (Sodium)

OpenStudy (anonymous):

More specifically Sodium-24

OpenStudy (anonymous):

\[Moles= \frac{mass_{grams}}{A_r}\] \[Atoms= Moles* Avogadro's-number\]

OpenStudy (anonymous):

A.N.= 6.0221415 *10^23

OpenStudy (anonymous):

I thought about that, but in my questions, avogadros number isn't mentioned until the next part (therefore, due to my experience with exam paper questions, it isn't needed until then)

OpenStudy (anonymous):

So initially the mass (N_0)\[=\frac{5*10^{-6}}{24}6.0221415 *10^{23}\]

OpenStudy (anonymous):

I'm not sure if there's any other (fundamental- i.e. not derived from it) way of working it out. Is this Homework or an exam paper?

OpenStudy (anonymous):

both, set an exam question for HW

OpenStudy (anonymous):

I'll try working it out with the arogarrdo number (or whatever it is called lol)

OpenStudy (anonymous):

Anyway, \[\huge N=125461281250000000*2^{-\frac{t_{hours}}{234}}\]

OpenStudy (anonymous):

Sorry, you're right, you don't need AN

OpenStudy (anonymous):

You just need\[\large 2^{-\frac{t}{234}}=\frac{1}{5}\]

OpenStudy (anonymous):

\[t=-234(\log_2(0.2))\]

OpenStudy (anonymous):

I got t = 1955669

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

=543.33 hours

OpenStudy (anonymous):

is my answer right?

OpenStudy (anonymous):

Its in seconds

OpenStudy (anonymous):

I got 1955992.2

OpenStudy (anonymous):

close enough (rounding errors)

OpenStudy (anonymous):

So yes. Did you do it using the equation 8 posts ago?

OpenStudy (anonymous):

*9

OpenStudy (anonymous):

no, I did this...

OpenStudy (anonymous):